OpenStudy (anonymous):

Find each percent of change and round to the nearest tenth. 1) 10 to 25 2) 70 to 82

OpenStudy (mertsj):

Find the change and divide by the original amount. 1. 25-10 = 15 so the change is 15 15/10 = 1.5 = 150% Can you do number 2?

OpenStudy (anonymous):

1) 82 / 70 = 1.1714 X 100 = 117.1% for number two?

hero (hero):

In general, to find the percent change from one number to another number, use the following formula: \[\% \Delta = \frac{|a-b|}{a} \times 100 \] a = original number b = new number

OpenStudy (mertsj):

You're getting warm. First you need to find the change. It it changes from 70 to 82, how much change is that?

OpenStudy (anonymous):

12? i thinkkkk. i'm not good at this.. lol.

hero (hero):

What about my formula? Do you understand it?

OpenStudy (anonymous):

yes,a little. its just easier for me when people write it out.

hero (hero):

So if 70 = a and 82 = b, you can't plug that into my formula and figure out what the percentage is?